A 25-foot ladder is placed against a building. The bottom of the ladder is 7 feet from the building. If the top of the ladder slips down 4 feet, how many feet will the bottom slide out? (It is not 4 feet).
7 squared+x squared=25 squared
49+x squared=625
x squared=576
x=24
Then what?

Answer by nerdybill(5411)

A 25-foot ladder is placed against a building. The bottom of the ladder is 7 feet from the building. If the top of the ladder slips down 4 feet, how many feet will the bottom slide out? (It is not 4 feet).
.
What you had is before the slip:
Let x = height where the ladder hits the building
7 squared+x squared=25 squared
49+x squared=625
x squared=576
x=24 feet
.
NOW, if it slips 4 feet -- that 24-4=20 feet
.
Let x = distance from wall:
x^2 + 20^2 = 25^2
x^2 + 400 = 625
x^2 = 225
x = 15 feet
.
But wait, the problem asked: "how many feet will the bottom slide out?"
Since, originally it was 7 feet from the wall
and now it is 15 feet
.
The answer is 8 feet.